OpenStudy (anonymous):

log x- log (x-2)= 4 both logs have a base of 2 find x, show steps

OpenStudy (anonymous):

the answer is 32/15 but i keep getting no solution

OpenStudy (misty1212):

\[\log_2\left(\frac{x}{x-2}\right)=4\] is how you want to begin

OpenStudy (solomonzelman):

\(\large\color{teal}{ \log_2x-\log_2(x-2)=4 }\) \(\large\color{teal}{ \log_2x=4+\log_2(x-2) }\) \(\large\color{teal}{ \log_2x=4\times \log_22+\log_2(x-2) }\) \(\large\color{teal}{ \log_2x= \log_22^4+\log_2(x-2) }\) continue....

OpenStudy (misty1212):

then write is in equivalent exponential form

OpenStudy (misty1212):

\[\frac{x}{x-2}=2^4\] is the exponential form

OpenStudy (misty1212):

or better yet \[\frac{x}{x-2}=16\]

OpenStudy (misty1212):

solve that for \(x\)

OpenStudy (misty1212):

you good with that ? that is how you are going to get the \(\frac{32}{15}\) that you know is the answer

OpenStudy (anonymous):

yeah, thanks